what is the center of mass for this object?

Find the center of mass of the following object.
Answer #1

here ,

let the centre of mass is Xcom , Ycom

as the centre of mass x - corrdinate

Xcom = (x1 * A1 + x2 * A2 + ..)/(total area)

Xcom = (10 * 10 * 5 + 10 * 20 * 15 + 15 * 5 * 22.5 + 5 * 10 * 30)/(10 * 10 + 10 * 20 + 15 * 5 + 10 * 5)

Xcom = 15.73 cm

for the y coordinate

Ycom = (y1 * A1 + y2 * A2 + ..)/(total area)

Ycom = (10 * 10 * 5 + 10 * 20 * 10 + 15 * 5 * 7.5 + 5 * 10 * 2.5)/(10 * 10 + 10 * 20 + 15 * 5 + 10 * 5)

Ycom = 7.5 cm

the centre of mass is (15.73 , 7.5) cm